Hello, I was wondering if you could help me out with this following problems. I have tried them and I am a few issues.
1) x^2 - 3x -10 >(underlined)0
2) X + 1 < 0
over
X - 5
3) 3x - 5 >(underlined)2
over
x + 2

Answer by stanbon(75887)

1) x^2 - 3x -10 >= 0
(x-5)(x+2) >= 0
Draw a number line and plot x = -2 and x = 5
---
Test a value in each of the 3 number-line intervals:
(x-5)(x+2) >= 0
Test x = -10: -*- > 0 ; true so solutions in (-oo,-5]
Test x = 0 : -*+ > 0 ; false
Test x = 10 : +*+ > 0 ; true so solutions in [2,+oo]
================================
2) (X + 1)/(x-5) < 0
Draw a number line and plot x = -1 and x = 5
Test a value in each of the 3 number-line intervals:
Test x = -10: -/- < 0 ; false
Test x = 0 : +/- < 0 ; true so solutions in [-1,5)
Test x = 10: +/+ < 0 ; false


==================================
3) (3x - 5)/(x+2) >= 2
Draw a number line and plot x = -2 and x = 5/3
Test as done in #1 and #2
=============================
Cheers,
Stan H.
=================================